OpenStudy (anonymous):

limit of e^tanx as x approaches pi/2 from the right I took the limit of e and multiplied it by pi/2 but cant get the answer. can someone explain please

OpenStudy (misty1212):

oh no

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first compute \[\lim_{x\to \frac{\pi}{2}^+}\tan(x)\]

OpenStudy (anonymous):

that would be positive infitiy right

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

you are approaching from the right

OpenStudy (anonymous):

negatiuve I mean sorry lol

OpenStudy (anonymous):

yeah so \(e^{-\infty}\) is the form

OpenStudy (anonymous):

okay that makes sense but now could you please explain to me why we take the limit of just the power? is it for any one with an exponent?

OpenStudy (anonymous):

It's a property of continuous functions. \(e^x\) is continuous for all values of \(x\), which allows you to say \[\large \lim_{x\to c}e^{f(x)}=e^{\lim\limits_{x\to c}f(x)}\]